1. A force 16 N pushes the 2 blocks (m1 = 3kg and m2 = 5 kg) in a horizontal surface but doesn’t move. What is the static coefficient friction of the blocks?



1
Expert's answer
2022-03-01T15:18:42-0500

By definition, the static friction force is equal to the pushing force "F_{fr} = F = 16N". On the other hand,



"F_{fr} = \mu N"

where "\mu" is the static friction coefficient, and "N = mg = (3kg+ 5kg)\cdot 9.8 = 78.4N" is the normal reaction. Thus, obtain:



"\mu = \dfrac{F}{N} = \dfrac{16N}{78.4N} \approx 0.2"

Answer. 0.2.